# The radiation condition for Helmholtz equations above locally perturbed periodic surfaces

### v1

@@ 4 The main theorems and the sketch of proof
To make the paper more readable, we here present the main theorems as well as the sketch of proof for the theorem. We will prove in the following results. First define the radiation condition as follows.
Definition 4.1.
@@
    the restriction $u\big{|}_{\mathbb{R}^{+}_{H+h}}$ satisfies the half-space Sommerfeld radiation condition, i.e.,
@@
The solution $u\in H^{1}_{loc}(\Omega)$ to (2.1) - (2.2) also satisfies the radiation condition defined in Definition 4.1.
@@
To treat these two different parts, we begin with the inverse Floqut-Bloch transform to obtain the solution $u$ to the field $w$:
$u(x+(2\pi j,0)^{\top})=\int_{\Lambda}w(\alpha,x)e^{\mathrm{i}2\pi\alpha}\,\mathrm{d}\alpha\,,\quad x\in\Omega^{H+h}_{0}.$ (4.3)
Note that $w$ satisfies either (3.7) or (3.8) holds. Based on these formulas, $u$ is decomposed into integrals in terms of the following forms with $\alpha_{0}\in\mathcal{S}$:
@@
We also note that (4.3) only describes the function in $\Omega^{H}$. For the definition in +$\mathbb{R}^{H}_{+}$, we need the apply the UPRC:
@@
Since $u$ is decomposed into the functions in terms of $u_{0}$, $u_{1}$ (or also $u_{2}$), we need to study the radiation condition for
@@
where $j=0,1,2$. We will need to study the decay rate of these three functions, as well as the radiation conditions for these functions, when $r=|x|\rightarrow\infty$.
To prove Theorem 4.2, we need to analyze the asymptotic behaviours of a family of integrals (see Table 5). This is a preparation step for the main proof, and will be given in Section 5.
For the proof for Part (A), we simply study the asymptotic behaviour of the terms (4.4), (4.5) and (4.6) in Section 6, based on the analysis of some integral in Section 5. We will prove that when $r\rightarrow\infty$, in the weak sense,
@@
The proof for Part (B) is the major part in this paper, where we prove the radiation conditions for $v_{0}$, $v_{1}$ and $v_{2}$. The proof for $v_{0}$ simply follows the proof in [20], using the decay rate of $u_{0}$ given by Section 6.1. We refer to Section 7 for the details. The proof for $v_{1}$ and $v_{2}$ is more complex, since the decay rates of $u_{1}$ and $u_{2}$ are too low to prove the theorem.
@@
$u_{1}(y_{1},H)=\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}w_{1}(\alpha,y)\,\mathrm{d}\alpha\,=\sum_{j\in\mathbb{Z}}\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}g_{j}(\alpha)e^{\mathrm{i}(\alpha+j)y_{1}}\,\mathrm{d}\alpha\,.$
Define
@@
$u_{1}(y_{1},H)=\int_{\mathbb{R}}h(\alpha)e^{\mathrm{i}\alpha y_{1}}\,\mathrm{d}\alpha\,\quad\Rightarrow\quad\mathcal{F}[\phi](\xi)=h(\xi).$
On the other hand, when we fix $x_{2}\geq H+h$,
@@
Thus from the convolution theorem, since
$v_{1}(x)=\int_{\mathbb{R}}\widetilde{\Phi}(x_{1}-y_{1})u_{1}(y_{1},H)\,\mathrm{d}y\,_{1}=\left(\widetilde{\Phi}*\phi\right)(x_{1}),$
the Fourier transform of $u$ has the form of
@@
Thus we need to study the radiation conditions for $v_{1}^{(j)}$ and $v_{2}^{(j)}$, respectively. For the investigation for $v_{1}^{(j)}$ we refer to Section 8 and for $v_{2}^{(j)}$ we refer to 9. Note that with different choices of $j$, the signs of $k^{2}-(\alpha+j)^{2}$ are also different. Thus we study the cases
@@
  • when $j\in J_{+}(\alpha_{0})$ in Sectin 8.2 for $v_{1}^{(j)}$ and in Section 9.2 for $v_{2}^{(j)}$;
  • when $j\in J_{0}(\alpha_{0})$ in Sectin 8.3 for $v_{1}^{(j)}$ and in Section 9.3 for $v_{2}^{(j)}$.
In Section 10, we finish the proof for Part (B) by summarizing the results from Section 7, 8 and 9. For the extension to locally perturbed periodic surfaces, we refer to Section 11.
@@
5 Analysis of some important integrals
To prepare for the estimation of the radiation condition in Theorem 4.2 with the help of the Floquet-Bloch transform, we need to estimate the decay of some integrals depends on a parameter, which are concluded in Table 5.
class name definition decay rate
We will consider the decay rate of the above integrals when $r\rightarrow\infty$. Here $\delta>0$ is a fixed small value. For simplicity, we all require that $\phi\in C^{\infty}_{0}(\mathbb{R})$ and vanishes either near $\delta$ (for example in $I^{A}_{1}(r),\,I^{A}_{2}(r),\,I^{B}_{1}(r),\,I^{B}_{2}(r)$) or the edge points $\gamma_{1},\gamma_{2}$ (for example in $I^{C}_{1}(r),\,I^{C}_{2}(r),\,I^{D}_{1}(r),\,I^{D}_{2}(r)$. Note that the results can be naturally extended to more general functions, i.e., the condition $\phi\in C^{\infty}_{0}(\mathbb{R})$ is always too strong (for example see Remark 5.2). We will not discuss the very explicit explanations since it will be too long for the readers. For the discussions of the integrals of Classes A, B, C and D, we refer to Subsections 5.1, 5.2, 5.3 and 5.4, respectively.
5.1 Analyse of the integrals of Class A
We begin with the first integral of Class A:
$I_{1}^{A}(r):=\int_{0}^{\delta}\frac{1}{\sqrt{x}}\phi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,,\quad r>0.$ (5.1)
Theorem 5.1.
$|I^{A}_{1}(r)|\leq Cr^{-1/2},\quad\forall\,r>0.$ (5.2)
Proof.
From direct computation,
where
where $S(x)$ and $C(t)$ are Fresnel integrals (see equation 11.7) and uniformly bounded. Thus
For (II), with integration by parts,
where
Since $\frac{1}{\sqrt{x}}$ is integrable in $[0,\delta]$
$\left|(II)\right|\leq\frac{C}{r}.$
Together with (I), we get that
$|I_{1}^{A}(r)|\leq C\frac{1}{\sqrt{r}}.$
Remark 5.2.
From the proof of Theorem 5.1, the positive constant $C$ only depends on the norm $\|\phi\|_{C^{2}[0,\delta]}$ thus the result is extended naturally for all $\phi\in C^{2}_{0}[0,\delta]$. However, for the rest of estimations in this section, we will not give very explicit discussions to save the space.
With the analysis of the first integral of Class A 5.2, we are prepared to analyze the second integral of Class A
$I_{2}^{A}(r):=\int_{0}^{\delta}x^{-1/2+m}\phi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,,\quad m\in\mathbb{N}.$ (5.3)
The estimation will be concluded as the corollary of Theorem 5.1.
Corollary 5.3.
$\left|I^{A}_{2}(r)\right|\leq Cr^{-1/2-m},\quad\forall\,r>0.$ (5.4)
Proof.
We prove by induction. For $m=0$, the result has already been proved in Theorem 5.1. Then assume that for a non-negative integer $m$, the estimation 5.4 holds. Then for $m+1$, from integration by parts, since $\phi(\delta)=0$,
$\int_{0}^{\delta}{x}^{-1/2+m+1}\phi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,$ $=\frac{1}{\mathrm{i}r}\int_{0}^{\delta}{x}^{-1/2+m+1}\psi(x)\,\mathrm{d}e\,^{\mathrm{i}rx}$
$=-\frac{1}{\mathrm{i}r}\int_{0}^{\delta}\left((-1/2+m+1)x^{-1/2+m}\phi(x)+x^{-1/2+m+1}\phi^{\prime}(x)\right)e^{\mathrm{i}rx}\,\mathrm{d}x\,$
$:=-\frac{1}{\mathrm{i}r}\int_{0}^{\delta}x^{-1/2+m}\psi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,$
where
which vanishes near $x=\delta$. With (5.2), there is a constant $C>0$ which only depends on $\psi$ (equivalently $\phi$) such that
$\left|\int_{0}^{\delta}{x}^{-1/2+m+1}\psi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,\right|\leq Cr^{-1}r^{-1/2-m}=Cr^{-1/2-m-1}.$
The proof is then finished. ∎
Of course we are able to estimate the asymptotic behaviours of the functions with smoother integrand, i.e., the third integral of Class A:
$I^{A}_{3}(r):=\int_{0}^{\delta}x^{m}\phi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,,\quad r>0,$ (5.5)
where $m=0,1,2,\dots$.
Theorem 5.4.
$\left|I^{A}_{3}(r)\right|\leq Cr^{-1-n}.$ (5.6)
Proof.
We prove by induction. First when $m=0$, from integration by parts,
Then we can immediately get that there is a $C>0$ such that
Now assume that the estimation (5.5) holds for any $m\geq 1$, then for $m+1$,
$=-\frac{1}{\mathrm{i}r}\int_{0}^{\delta}\left[(m+1)x^{m}\phi(x)+x^{m+1}\phi^{\prime}(x)\right]e^{\mathrm{i}rx}\,\mathrm{d}x\,$
$:=-\frac{1}{\mathrm{i}r}\int_{0}^{\delta}x^{m}\psi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,,$
where
which also vanishes around $\delta$. With (5.6),
Thus (5.6) also holds for $m+1$. The proof is finished. ∎
5.2 Analyse of the integrals of Class B
The first integral of Classe B is defined as:
$I_{1}^{B}(r):=\int_{0}^{\delta}\sqrt{x}\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,,\quad r>0,$ (5.7)
where the function $\phi$ has exactly the same property as in Section 5.1.
Theorem 5.5.
$\left|I_{1}^{B}(r)\right|\leq Cr^{-3/4},\quad\forall\,r>0.$ (5.8)
Proof.
Similar to the proof of Theorem 5.1, we can also write
where
$\psi(x)=\frac{\phi(x)-\phi(0)}{x}\in C^{\infty}(\mathbb{R}).$
For (I), let $s:=\sqrt{r}x$, from (11.8)
For (II), with integration by parts,
where $C>0$ depends on $\psi$ and $\widetilde{\psi}$, which is defined by
So finally we can show that
$|I^{B}_{1}(r)|\leq Cr^{-3/4},\quad\forall\,r>0.$
Now we consider the second integral of Class B:
$I_{2}^{B}(r)=\int_{-a}^{\delta}(x+a)^{2}\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,.$ (5.9)
Theorem 5.6.
$\left|I^{B}_{2}(r)\right|\leq Cr^{-1/2}.$ (5.10)
Proof.
$e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,=\frac{1}{2\mathrm{i}rx}\,\mathrm{d}e\,^{\mathrm{i}rx^{2}}.$
To avoid the singularity at 00, we split the integral (5.9) into three parts:
Note that when $a<r^{-1/2}$, then (I) vanishes and the integral interval of (II) is modified as $[-a,r^{-1/2}]$. For simplicity, let ${\varepsilon}_{0}=\min\{a,r^{-1/2}\}$ then ${\varepsilon}\leq r^{-1/2}$. For (II), from direct computation,
Note that here
$|I^{B}_{2}(r)|\leq Cr^{-1/2}.$
5.3 Analyse of the integrals of Class C
$I_{1}^{C}(r)=\int_{\gamma_{1}}^{\gamma_{2}}\frac{1}{\sqrt{x-a}}\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,;$ (5.11)
[…]
Theorem 5.7.
Suppose $\phi\in C_{0}^{\infty}(\gamma_{1},\gamma_{2})$, where $\gamma_{1}<0<a<\gamma_{2}$. Then there is a constant which only depends on $\phi$ such that
$\left|I_{1}^{C}(r)\right|\leq Ca^{-1}r^{-1/2}.$ (5.13)
Moreover, for any $|a|<1$, we have
$\left|a\,I_{1}^{C}(r)\right|\leq Cr^{-1/2}.$ (5.14)
Proof.
Note that similarly to the proof of Theorem 5.6, the point 00 becomes a singularity when we apply integration by parts to the term $e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,$. From the definition (5.11), $a$ is another singularity. To avoid these singularities, We split the integral $I_{1}^{C}(r)$ into five or three parts, according to different situations. When $a\geq 2r^{-1/2}$, then $a-1/r>2r^{-1/2}-r^{-1/2}=r^{-1/2}$. Then
$:=(I_{1})+(II_{1})+(I_{2})+(II_{2})+(I_{3}).$ (5.15)
$I_{1}^{C}(r)=\int_{\gamma_{1}}^{-r^{-1/2}}+\int_{-r^{-1/2}}^{4r^{-1/2}}+\int_{4r^{-1/2}}^{\gamma_{2}}\frac{1}{\sqrt{x-a}}\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,:=(I_{1})+(II_{3})+(I_{4}).$ (5.16)
Similarly, $(II_{3})$ involves the singularities.
(i) Case I: $a\geq 2r^{-1/2}$, where $I^{C}_{1}(r)$ is given by (5.15). We need to estimate two terms $(II_{1})$ and $(II_{2})$ which involve singularities, and the rest of the three terms which don’t.
First observe that since $a\geq 2r^{-1/2}$, from direct computation,
and
[…]
where
and
Since
$\left|(I_{\ell})\right|\leq\frac{1}{2r}\left|\left.\frac{1}{x\sqrt{x-a}}\phi(x)\right|_{+}\right|+\frac{1}{2r}\left|\left.\frac{1}{x\sqrt{x-a}}\phi(x)\right|_{-}\right|+\frac{1}{2r}|(i_{\ell})|,$
we need to estimate the values $\left|\frac{1}{x\sqrt{x-a}}\phi(x)\right|$ at the boundaries of $K_{\ell}$, $\ell=1,2,3$, as well as the integral $|(i_{\ell})|$. Note that since $\phi(\gamma_{1})=\phi(\gamma_{2})=0$, we only need to consider the points $x=\pm r^{-1/2},\,a\pm r^{-1}$. From direct computation,
$\left|\left.\frac{1}{x\sqrt{x-a}}\right|_{x=-r^{-1/2}}\right|=\frac{\sqrt{r}}{\sqrt{a+r^{-1/2}}}\leq\sqrt{\frac{r}{a}};$
since $a\geq 2r^{-1/2}$, $a-r^{-1/2}\geq a/2$ thus
$\left|\left.\frac{1}{x\sqrt{x-a}}\right|_{x=r^{-1/2}}\right|=\frac{\sqrt{r}}{\sqrt{a-r^{-1/2}}}\leq C\sqrt{\frac{r}{a}};$
and since $a\pm r^{-1}\geq a-r^{-1}\geq a-a^{2}/4>a/2$,
Then we only need to estimate the integral $(i_{\ell})$. We need to discuss for each interval seperately. For the integral on $K_{1}=[\gamma_{1},-r^{-1/2}]$,
For $K_{3}=[a+1/r,\gamma_{2}]$,
Thus altoghter, there is a constant $C>0$ such that
$|(i_{2})|\leq C\frac{\sqrt{r}}{a}$
when $a\geq 2r^{-1/2}$. Thus combine with previous estimations we have
The first case is then proved.
(ii) Case II: $0\leq a<2r^{-1/2}$, when $I^{C}_{1}(r)$ is given by (5.16). Similarly, we need to consider the term $(II_{3})$ which involves the singularities, and $(I_{1})$ and $(I_{4})$ which are regular. Notice that for $(I_{1})$, it has been discussed in Case I and the estimation still holds for $a<2r^{-1/2}$, thus we don’t need to discuss here. Thus we only need to estimate the terms $(II_{3})$ and $(I_{4})$.
For $(II_{3})$, from direct computation,
$=2\sqrt{a+r^{-1/2}}+2\sqrt{4r^{-1/2}-a}\leq Cr^{-1/4}.$
Note that since $a<2r^{-1/2}$, we estimate that $r^{1/4}<\sqrt{{2}}a^{-1/2}$ then
$|(II_{3})|\leq Cr^{-1/2}r^{1/4}\leq Ca^{-1/2}r^{-1/2}.$
To estimate $(I_{4})$, we define the following interval similary
$K_{4}=[4r^{-1/2},\gamma_{2}].$
With integration by parts, since $\phi(\gamma_{2})=0$,
[…]
From direct computation, since $a<2r^{-1/2}$, $4r^{-1/2}-a>a$ then
Now we move on to the integral $(i_{4})$. Similar to Case I,
Similar to Case I, we can also estimate that
$\left|I^{C}_{1}(r)\right|\leq C\frac{1}{a\sqrt{r}}$
holds uniformly for $r$ and $10<a<$ 1. Then we can get the estimation (5.14) immediately. ∎
As the corollary from Theorem 5.7, we estimate the second integral of Class C, i.e., $I^{C}_{2}(r)$.
Corollary 5.8.
$\left|I_{2}^{C}(r)\right|\leq Cr^{-1/2}.$ (5.17)
Proof.
From direct computation,
[…]
From (5.14), $|aI_{1}^{C}(r)|\leq Cr^{-1/2}$, we only need to consider the integral (I). First split the integral into three parts due to the singularity $a$:
$(I)=\int_{\gamma_{1}}^{a-1/r}+\int_{a-1/r}^{a+1/r}+\int_{a+1/r}^{\gamma_{2}}\frac{1}{\sqrt{x-a}}\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,:=(i_{1})+(ii)+(i_{2}).$
From direct computation,
For $(i_{1})$, with integration by parts,
$(i_{1})$ $=\int_{\gamma_{1}}^{a-1/r}\frac{1}{\sqrt{x-a}}\phi(x)\,\mathrm{d}e\,^{\mathrm{i}rx^{2}}$
where
Note that
and
Thus we can easily obtain that
$|(i_{1})|\leq Cr^{-1/2}.$
With similar technique, we can also get
$|(i_{2})|\leq Cr^{-1/2}.$
Together with the estimation of (ii),
$|(I)|\leq Cr^{-1/2}.$
The proof is then finished. ∎
5.4 Analyse of the integrals of Class D
The first integral of Class D is defined as
$I^{D}_{1}(r):=\int_{\gamma_{1}}^{\gamma_{2}}|x-a|\phi(x)e^{\mathrm{i}rx^{2}}\,\mathrm{d}x\,$ (5.18)
Theorem 5.9.
$\left|I^{D}_{1}(r)\right|\leq Cr^{-1/2}.$ (5.19)
Proof.
We only need to consider the case that $a\in(\gamma_{1},\gamma_{2})$. When $a\not\in(\gamma_{1},\gamma_{2})$, we can easily apply the results in Theorem 5.4 to obtain the expected results. Thus the details are omitted.
$:=a(I)-a(II)-(III)+(IV).$
where
For $(I_{1})$, let $s=\sqrt{2r}t$ with the Fresnel integral (11.7),
For $(I_{2})$, with integration by parts,
Thus we can get the estimations for both $(I)$ and $(II)$:
$|(I)|,\,|(II)|\leq Cr^{-1/2}.$
$|(III)|,\,|(IV)|\leq Cr^{-1}.$
Now we move on to the second integral of Class D, which is defined by
$I^{D}_{2}(r):=\int_{\gamma_{1}}^{\gamma_{2}}{\rm sgn}(x-a)\phi(x)e^{\mathrm{i}rt^{2}}\,\mathrm{d}t\,.$ (5.20)
Note that here the sign function ${\rm sgn}$ is defined as
The decay of $I^{D}_{2}(r)$ is concluded as the corollary of Theorem 5.7.
Corollary 5.10.
$\left|I^{D}_{2}(r)\right|\leq Cr^{-1/2}.$ (5.21)
Proof.
Directly from the definition of the sign function,
where both functions has exactly the same forms of the terms (I) and (II) in the proof of Theorem 5.7. Thus the proof follows exactly as the proof of the previous theorem. ∎

### v2

@@ 4 The main theorems and the sketch of proof
To make the paper more readable, we here present the main theorem as well as the sketch of proof in this section. First define the radiation condition as follows.
Definition 4.1 (Radiation Condition).
@@
    the restriction $u\big{|}_{U_{H+h}}$ satisfies the half-space Sommerfeld radiation condition, i.e.,
@@
Let $\Gamma$ be the graph of a periodic bounded function. The unique solution $u\in H^{1}_{r}(\Omega)$ ($|r|<1$) to (2.1) - (2.2) also satisfies the radiation condition defined in Definition 4.1.
@@
To prove Theorem 4.2, we begin with the representation (3.8) and (3.10) with different $k$’s. With the $\alpha$-quasi-periodicities of the functions, the solution $u$ is decomposed into integrals in the following three forms:
@@
We also note that (4.3), (4.4) and (4.5) only describe the solutions restricted in $\Omega^{H+h}$. For the extension of the solution into the half space $U_{H}$, we need the apply the UPRC (3.8)
@@
Since $u$ is decomposed into the functions in terms of $u_{0}$, $u_{1}$ and $u_{2}$, we need to study the radiation condition for
@@
where $j=0,1,2$. We will need to study the decay rate of these three functions, as well as the radiation conditions for these functions, when $r=|x-(0,H)^{\top}|\rightarrow\infty$.
For the proof for Part (A), we simply study the asymptotic behaviour of the terms (4.3), (4.4) and (4.5) in Section 5, based on the analysis of some integral in Section B. We will prove that when $r\rightarrow\infty$, in the weak sense,
@@
The proof for Part (B) is the more challenging part, where we prove the radiation conditions for $v_{0}$, $v_{1}$ and $v_{2}$. The proof for $v_{0}$ simply follows the proof in [19], using the decay rate of $u_{0}$ given by Section 5.1. We refer to Section 6 for the details. The proofs for $v_{1}$ and $v_{2}$ are more complex. Since the decay rates of $u_{1}$ and $u_{2}$ are too slow, the method for $v_{0}$ no longer works thus new approaches will be introduced.
@@
$u_{1}(y_{1},H)=\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}w_{1}(\alpha,y)\,\mathrm{d}\alpha\,=\sum_{j\in\mathbb{Z}}\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}g_{j}(\alpha)e^{\mathrm{i}(\alpha+j)y_{1}}\,\mathrm{d}\alpha\,,$
where the interchanging of the summation and integration is guaranteed by the decay of $g_{j}(\alpha)$. Define
@@
$u_{1}(y_{1},H)=\int_{\mathbb{R}}h(\alpha)e^{\mathrm{i}\alpha y_{1}}\,\mathrm{d}\alpha\,\quad\Rightarrow\quad\mathcal{F}[\phi](\xi)=h(\xi)$
where $\mathcal{F}$ is the Fourier transform. On the other hand, when we fix $x_{2}\geq H+h$,
@@
Since
$v_{1}(x)=\int_{\mathbb{R}}\widetilde{\Phi}(x_{1}-y_{1})u_{1}(y_{1},H)\,\mathrm{d}y\,_{1}=\left(\widetilde{\Phi}*\phi\right)(x_{1})$
where * is the convolution operator. From the convolution theorem, the Fourier transform of $u$ has the form of
@@
Thus we need to study the radiation conditions for $v_{1}^{(j)}$ and $v_{2}^{(j)}$, respectively. For the investigation for $v_{1}^{(j)}$ we refer to Section 7 and for $v_{2}^{(j)}$ we refer to 8. Note that with different choices of $j$, the signs of $k^{2}-(\alpha+j)^{2}$ are also different. Thus we study the cases
@@
  • when $j\in J_{+}(\alpha_{0})$ in Sectin 7.2 for $v_{1}^{(j)}$ and in Section 8.2 for $v_{2}^{(j)}$;
  • when $j\in J_{0}(\alpha_{0})$ in Sectin 7.3 for $v_{1}^{(j)}$ and in Section 8.3 for $v_{2}^{(j)}$.
In Section 9, we finish the proof for Part (B) by summarizing the results from Section 6, 7 and 8. For the extension to locally perturbed periodic surfaces, we refer to Section 10.